OpenStudy (anonymous):

Can i get help finding three consecutive even integers such that when the first integer is multiplied by the third integer, the result is 2 more than 5 times the second integer.

OpenStudy (raden):

let that number are, x, x+1, x+2 according information above : x(x+2) = 5(x+1) + 2 solve for x

OpenStudy (anonymous):

i see. but i got stuck when i multiplied x and x because it equals x^2

OpenStudy (anonymous):

i dont know what to do with it when i add like terms

OpenStudy (raden):

simplify this, first x(x+2) = 5(x+1) + 2 x^2+2x = 5x+5+2 x^2+2x = 5x+7 x^2+2x-5x-7=0 x^2-3x-7=0

OpenStudy (raden):

but, look no x's are integer number there !!

OpenStudy (anonymous):

do i solve for x in a algebraic solution?

OpenStudy (raden):

yes, use the quadratic formula

OpenStudy (raden):

quadratic equation formula i meant

OpenStudy (anonymous):

isnt that -b/2(a)?

OpenStudy (raden):

that is the sum of roots formula

OpenStudy (raden):

oppps... that is the formula to get x-vertex :) the sum of roots is -b/a, sorry

OpenStudy (raden):

hai, i come back :)

OpenStudy (raden):

well, i want to show u how to get x's use quadratic equation formula :